OpenStudy (anonymous):

Can any one please help me!!!!!!!!!!!! Solution A is 40% acid and solution B is 80% acid. How much of each should be used to make 300 liters of a solution that is 64% acid? A. Solution A is 150 liters and Solution B is 150 liters. B. Solution A is 100 liters and Solution B is 200 liters. C. Solution A is 120 liters and Solution B is 180 liters.

OpenStudy (neer2890):

check out the options: A: if A is 150 liters then %age of acid =(40/100)*150= 60 litres if B is 150 liters then %age of acid =(80/100)*150= 120 litres In total mixture, total quantity of acid= 60+120=180 litres and we know that 64% of 300 =192 but we have 180 liters. so option A is wrong.

OpenStudy (neer2890):

for option C: Quantity of acid in A=0.40*120=48 liters Quantity of acid in B=.80*180= 144 liters so in mixture, Quantity of acid = 144+48=192 liters which is 64% of 300 liters. so C is correct.